two numbers have a sum of 24.one number in 6 less than twice,the other.what are the numbers?
Answer by ptaylor(2198)

Let x=the other number
Then 2x-6 =the first number
Now we are told that:
x+2x-6=24 add 6 to both sides
x+2x-6+6=24+6 collect like terms
3x=30 divide both sides by 3
x=10---------------------------------the other number
2x-6=2*10-6=14-----------------------------the first number
CK
10+14=24
24=24
also 2*10-6=14
14=14
Hope this helps---ptaylor